# Towards Efficient Training of Graph Neural Networks: A Multiscale Approach

### v2

@@ 6.1 Transductive Learning
Random and subgraphs(p=1) $2.67\times 10^{2}$ 79.9% (1.4%) $8.10\times 10^{2}$ 71.2% (0.8%) $7.36\times 10^{2}$ 76.8% (0.5%)
Random and subgraphs(p=2) $3.46\times 10^{2}$ 80.2% (1.1%) $8.57\times 10^{2}$ 70.8% (0.4%) $9.87\times 10^{2}$ 77.5% (1.1%)
Table 2: Comparison of different methods for Cora, CiteSeer, and PubMed datasets. We present the test accuracy for multiscale training with 3 levels, as well as FLOPs analysis, in $10^{6}$ units. The cheapest FLOPs count for each network and dataset appears in red. Results for fine grid appear in blue.
We explore two pooling strategies—random and Top-k—along with subgraph generation via $k$-hop neighborhoods and a hybrid approach combining both. Each method is tested under two connectivity settings: the original adjacency ($p=1$) and an enhanced version ($p=2$), where we square the adjacency matrix. We report test accuracy across three runs with different random seeds and include standard deviation to assess robustness. Our results indicate that enhanced connectivity improves accuracy in some settings. Table 2 further presents three levels of multiscale training alongside theoretical FLOP analyses, demonstrating the efficiency of our approach. The details of the FLOPs analysis are provided in 5, and additional results for two and four levels of multiscale training are provided in Appendix A.
2 levels 3 levels 4 levels
Model Method FLOPs Time Test Acc FLOPs Time Test Acc FLOPs Time Test Acc
GCN Fine grid          1 level: $1.96\times 10^{4}$ 39.70 71.76% (0.1%)
Random and subgraphs $5.49\times 10^{3}$ 37.44 70.71% (0.2%) $2.70\times 10^{3}$ 34.50 70.93% (0.0%) $1.34\times 10^{3}$ 33.77 70.53% (0.2%)
GIN Fine grid          1 level: $3.75\times 10^{4}$ 61.66 68.39% (0.6%)
Random and subgraphs $1.06\times 10^{4}$ 39.56 65.87% (6.6%) $5.25\times 10^{3}$ 31.16 64.79% (3.0%) $2.62\times 10^{3}$ 32.36 67.46% (1.2%)
GAT Fine grid          1 level: $1.13\times 10^{4}$ 54.71 71.07% (0.2%)
Random and subgraphs $5.18\times 10^{3}$ 43.05 70.09% (0.2%) $2.56\times 10^{3}$ 35.79 69.62% (0.2%) $1.27\times 10^{3}$ 35.33 70.06% (0.2%)
Table 3: Comparison of different methods and coarsening levels for OGBN-Arxiv dataset. Time represents average epoch time in the observed levels of coarsening, in $ms$ units. The best time for a given network and level appear in orange. FLOP values represent computational cost at the achieved level, in $10^{6}$ units. The cheapest FLOPs count for each network and dataset appears in red. Results for fine grid appear in blue.
We evaluate our approach on large-scale graphs using the OGBN-Arxiv dataset Hu et al. (2020) with GCN Kipf & Welling (2016), GIN Xu et al. (2018), and GAT Velickovic et al. (2017). This dataset consists of 169,343 nodes and 1,166,243 edges. Using the same experimental setup as before, we present our results in Table 3, demonstrating that our multiscale framework achieves comparable or superior performance relative to standard training. Notably, random pooling proves to be the most effective approach, offering both efficiency and strong performance. Given the large number of edges in this dataset, we use $p=1$ and report the corresponding edge counts in Table 4. The results highlight the substantial reduction in edges at each pooling level, significantly improving training efficiency. In Table 3, we also provide a timing analysis, including both theoretical FLOP estimates and empirical runtime measurements, confirming that our multiscale approach enhances efficiency while maintaining strong predictive performance.
Additionally, consider the Flickr Zeng et al. (2019), WikiCS Mernyei & Cangea (2020), DBLP Bojchevski & Günnemanng (2017), the transductive versions of Facebook, BlogCatalog, and PPI Yang et al. (2020) datasets, detailed in Appendix B. We present results for Flickr and PPI, along with timing evaluations, as well as full results for all those datasets and additional experimental details, in Appendix B.
Method 1st Coarsening 2nd Coarsening 3rd Coarsening
No Coarsening          1 level: 1,166,243
Random 280,009 76,559 17,978
Topk 815,023 480,404 238,368
Table 4: Edge number for coarsening strategies for the coarse-to-fine method, with the OGBN-Arxiv dataset. Fine grid edge number is in blue .
6.2 Inductive Learning
We evaluate our method on the ShapeNet dataset Chang et al. (2015), a point cloud dataset with multiple categories, for a node segmentation task. Specifically, we present results on the ’Airplane’ category, which contains 2,690 point clouds, each consisting of 2,048 points. Nodes are classified into four categories: fuselage, wings, tail, or engine. For this dataset, we use the Dynamic Graph Convolutional Neural Network (DGCNN) Wang et al. (2018), which dynamically constructs graphs at each layer using the $k$-nearest neighbors algorithm, with $k$ as a tunable hyperparameter. We test our multiscale approach with random pooling, subgraph pooling, and a combination of both. In subgraph pooling, given that the data resides in $\mathbb{R}^{3}$, we construct subgraphs by selecting a random node and its $N_{j}$ nearest neighbors. Experiments are conducted using three random seeds, and we report the mean Intersection-over-Union (IoU) scores along with standard deviations in the top section of Table 5. Our results show that for all tested values of $k$, applying either random or subgraph pooling improves performance compared to standard training while reducing computational cost. Furthermore, combining both pooling techniques consistently yields the best results across connectivity settings, suggesting that this approach captures both high- and low-frequency components of the graph while maintaining training efficiency. Additional results for multiscale training on the full ShapeNet dataset are provided in Appendix C.
We apply the Multiscale Gradients Computation method to the ShapeNet ’Airplane’ category, evaluating various pooling techniques using Algorithm 2 with two levels. As shown in the bottom section of Table 5, our approach reduces training costs while achieving performance comparable to fine-grid training. These results show that the early stages of optimization can be efficiently performed using an approximate loss, as we propose, with fine-grid calculations needed only once the loss function nears its minimum. Additional results, timing analyses and experimental details for the ShapeNet dataset are provided in Appendix C.
Training using Multiscale Training
Fine Grid Random Subgraphs Random and Subgraphs
k=6 78.36% (0.28%) 79.86% (0.87%) 79.31% (0.20%) 80.01% (0.21%)
k=10 79.71% (0.23%) 80.41% (0.14%) 80.65% (0.15%) 80.97% (0.10%)
k=20 81.06% (0.08%) 81.37% (0.05%) 81.18% (0.32%) 81.89% (0.58%)
Training using Multiscale Gradient Computation Method
Fine Grid Random Subgraphs Random and Subgraphs
k=6 78.36% (0.28%) 79.42% (1.40%) 79.35% (0.12%) 79.62% (0.93%)
k=10 79.71% (0.23%) 80.00% (0.33%) 80.15% (0.71%) 80.19% (1.63%)
k=20 81.06% (0.08%) 81.30% (0.41%) 81.17% (0.25%) 81.12% (0.24%)
Table 5: Results on Shapenet ’Airplane’ dataset using DGCNN. Results represent Mean test IoU.
We further evaluate our Multiscale Gradients Computation method on the protein-protein interaction (PPI) dataset Zitnik & Leskovec (2017), consisting of graphs representing different human tissues. The results, with additional details, are provided in Appendix D.
7 Conclusions
Appendix A Additional Results for Planetoid Datasets
We provide additional results for Cora, Citeseer, and Pubmed datasets (Sen et al., 2008), for 2, 3, and 4 levels of multiscale training, using datasets details provided in 6.1. We test our methods using GCN (Kipf & Welling, 2016), GIN (Xu et al., 2018), and GAT (Velickovic et al., 2017). Results are presented in Table 6, where we show that across various levels of multiscale training and different methods, we achieve consistently better or comparable results to those of the fine grid original training while using a much cheaper training process.
Method $\downarrow$ / Dataset $\rightarrow$ Cora CiteSeer PubMed
2 levels 3 levels 4 levels 2 levels 3 levels 4 levels 2 levels 3 levels 4 levels
GCN Fine grid          1 level: 82.9% (1.1%)          1 level: 69.6% (1.0%)          1 level: 77.5% (0.4%)
Random(p=1) 81.7% (0.6%) 82.0% (0.5%) 81.7% (0.4%) 70.5% (0.3%) 70.4% (0.2%) 70.7% (0.2%) 77.6% (0.8%) 77.5% (0.6%) 77.5% (0.4%)
Random(p=2) 83.2% (0.6%) 83.0% (0.5%) 81.5% (0.3%) 66.6% (2.6%) 66.0% (0.8%) 66.2% (2.2%) 74.2% (0.4%) 73.8% (0.4%) 74.2% (0.7%)
Topk(p=1) 79.5% (0.7%) 79.9% (4.4%) 79.8% (2.2%) 68.5% (1.1%) 68.1% (0.1%) 67.4% (0.3%) 77.8% (0.9%) 78.2% (0.3%) 78.2% (0.6%)
Topk(p=2) 80.5% (2.5%) 79.6% (2.4%) 80.8% (0.8%) 66.4% (0.8%) 63.6% (2.7%) 65.8% (2.1%) 75.4% (0.4%) 74.5% (0.4%) 74.2% (0.6%)
Subgraphs(p=1) 82.7% (0.7%) 83.5% (0.4%) 81.3% (0.1%) 71.6% (0.5%) 71.2% (0.4%) 71.3% (0.4%) 77.7% (0.2%) 78.0% (0.3%) 78.7% (1.3%)
Subgraphs(p=2) 82.8% (0.4%) 83.1% (0.5%) 81.8% (0.7%) 71.5% (0.5%) 70.9% (0.2%) 70.9% (0.4%) 74.8% (1.2%) 76.1% (0.3%) 76.5% (1.1%)
Random and subgraphs(p=1) 81.7% (0.4%) 81.5% (0.3%) 81.5% (0.3%) 71.4% (0.5%) 70.0% (0.0%) 70.7% (0.3%) 77.4% (0.5%) 77.8% (0.8%) 76.7% (0.3%)
Random and subgraphs(p=2) 82.7% (0.3%) 82.6% (0.5%) 80.8% (0.1%) 70.8% (0.3%) 70.9% (0.2%) 70.7% (0.1%) 74.1% (0.7%) 73.9% (0.4%) 75.6% (0.9%)
GIN Fine grid          1 level: 76.4% (1.6%)          1 level: 63.8% (2.6%)          1 level: 75.4% (0.4%)
Random(p=1) 76.6% (1.5%) 76.2% (2.7%) 77.6% (2.2%) 66.3% (7.0%) 66.5% (2.2%) 66.4% (1.1%) 75.6% (1.7%) 75.7% (2.0%) 76.2% (1.3%)
Random(p=2) 73.1% (0.3%) 77.2% (0.6%) 71.6% (1.6%) 60.6% (2.6%) 60.5% (0.8%) 62.8% (1.3%) 75.6% (2.9%) 75.1% (0.9%) 74.3% (1.2%)
Topk(p=1) 77.5% (1.3%) 77.6% (5.6%) 79.4% (1.9%) 67.3% (1.3%) 64.8% (3.9%) 66.5% (1.5%) 76.6% (0.8%) 75.6% (1.6%) 75.5% (1.6%)
Topk(p=2) 76.2% (1.6%) 74.6% (3.0%) 75.6% (2.3%) 61.1% (1.1%) 60.5% (2.0%) 60.0% (2.9%) 76.0% (1.8%) 76.5% (1.5%) 66.1% (6.4%)
Subgraphs(p=1) 77.5% (0.9%) 78.7% (1.1%) 78.2% (1.4%) 68.2% (0.5%) 68.8% (1.6%) 68.7% (1.4%) 75.8% (1.2%) 77.2% (1.2%) 77.2% (0.7%)
Subgraphs(p=2) 76.2% (8.57%) 76.9% (1.72%) 77.3% (8.55%) 60.1% (0.29%) 65.0% (1.8%) 66.6% (1.0%) 75.6% (2.8%) 75.2% (1.5%) 55.4% (4.7%)
Random and subgraphs(p=1) 80.2% (2.5%) 76.9% (1.0%) 76.8% (2.1%) 66.0% (0.9%) 65.3% (1.2%) 65.5% (1.1%) 75.2% (0.8%) 76.1% (2.8%) 75.0% (0.7%)
Random and subgraphs(p=2) 75.5% (0.7%) 75.9% (0.8%) 74.6% (22.5%) 64.1% (2.2%) 66.0% (0.2%) 66.7% (1.0%) 77.8% (5.6%) 75.9% (0.5%) 76.4% (2.4%)
GAT Fine grid          1 level: 79.6% (2.05%)          1 level: 67.2% (6.5%)          1 level: 75.5% (0.5%)
Random(p=1) 80.1% (0.8%) 80.3% (0.2%) 79.8% (0.8%) 71.1% (0.4%) 72.9% (1.5%) 71.4% (0.3%) 76.5% (0.3%) 77.5% (1.5%) 76.4% (0.8%)
Random(p=2) 82.0% (0.5%) 79.5% (0.6%) 79.5% (1.9%) 71.6% (1.1%) 70.7% (1.0%) 70.6% (1.5%) 75.2% (1.0%) 77.4% (1.2%) 78.6% (2.5%)
Topk(p=1) 77.3% (1.8%) 79.8% (0.6%) 77.5% (2.9%) 67.2% (1.5%) 66.8% (0.6%) 70.8% (1.3%) 76.1% (0.2%) 76.1% (0.9%) 75.6% (2.5%)
Topk(p=2) 79.6% (2.3%) 76.8% (6.7%) 79.0% (1.6%) 67.4% (2.2%) 65.2% (0.7%) 67.0% (0.2%) 74.2% (1.0%) 74.9% (0.7%) 73.3% (1.0%)
Subgraphs(p=1) 80.2% (2.6%) 78.5% (0.8%) 79.5% (0.8%) 69.1% (0.6%) 70.2% (0.6%) 68.4% (2.7%) 77.0% (1.0%) 75.4% (4.6%) 75.8% (1.6%)
Subgraphs(p=2) 80.5% (1.5%) 80.7% (1.2%) 79.9% (0.9%) 70.1% (0.5%) 68.6% (0.3%) 66.0% (0.2%) 75.1% (0.85%) 77.3% (1.1%) 78.2% (3.0%)
Random and subgraphs(p=1) 80.6% (0.2%) 79.9% (1.4%) 80.3% (0.9%) 70.8% (0.1%) 71.2% (0.8%) 71.2% (0.3%) 77.3% (0.8%) 76.8% (0.5%) 76.0% (1.9%)
Random and subgraphs(p=2) 81.1% (1.1%) 80.2% (1.1%) 78.7% (0.4%) 70.3% (0.4%) 70.8% (0.4%) 71.1% (0.3%) 75.7% (1.8%) 77.5% (1.1%) 76.4% (1.9%)
Table 6: Comparison of different methods and coarsening levels for Cora, CiteSeer, and PubMed datasets.
Appendix B Additional Results for Transductive Learning
We further test our methods using additional datasets. We provide results for Flickr Zeng et al. (2019) and PPI (transductive) Yang et al. (2020), showing both timing analysis, in Table 7, and full multiscale training results in Table 8. Our results show that while achieving comparable accuracy to the original fine grid, our methods show significant timing improvement both in a theoretical manner as well as in FLOPs analysis, which we calculate as explained in Appendix 5.
@@ Appendix B Additional Results for Transductive Learning
FLOPs Time Test acc FLOPs Time Test acc
GCN Fine grid $1.72\times 10^{4}$ 68.90 54.77% (0.2%) $8.09\times 10^{3}$ 64.66 70.74% (0.2%)
@@
Random and subgraphs $4.44\times 10^{3}$ 49.74 55.85% (1.6%) $1.63\times 10^{3}$ 68.06 71.50% (0.0%)
GIN Fine grid $3.58\times 10^{4}$ 76.29 52.05% (1.0%) $1.34\times 10^{4}$ 48.76 85.67% (7.2%)
@@
Random and subgraphs $9.01\times 10^{3}$ 53.51 52.00% (1.8%) $3.16\times 10^{3}$ 68.52 90.83% (12.3%)
GAT Fine grid $1.53\times 10^{4}$ 103.54 52.71% (0.9%) $4.57\times 10^{3}$ 192.20 82.37% (3.1%)
@@
Random and subgraphs $3.99\times 10^{3}$ 62.38 53.26% (0.2%) $9.21\times 10^{2}$ 79.21 87.03% (0.8%)
Table 7: Comparison of different methods for Flickr and PPI (transductive) dataset. The test accuracy is calculated for 3 levels of multiscale training. Time represents average epoch time in the observed levels of coarsening, in $ms$ units. The best time for a given network appears in orange. FLOP values represent computational cost at the achieved level, in $10^{6}$ units. The cheapest FLOPs count for each network and dataset appears in red. Results for fine grid appear in blue.
@@
2 levels 3 levels 4 levels 2 levels 3 levels 4 levels
GCN Fine grid          1 level: 54.77% (0.2%)          1 level: 70.74% (0.2%)
@@
Random and subgraphs 55.65% (0.1%) 55.85% (1.6%) 55.08% (0.2%) 70.88% (0.1%) 71.50% (0.0%) 72.89% (0.1%)
GIN Fine grid          1 level: 52.05% (1.0%)          1 level: 85.67% (7.2%)
@@
Random and subgraphs 52.74% (1.0%) 52.00% (1.8%) 52.50% (2.0%) 84.25% (4.1%) 90.83% (12.3%) 93.31% (12.1%)
GAT Fine grid          1 level: 52.71% (0.9%)          1 level: 82.37% (3.1%)
@@
Random and subgraphs 52.70% (0.0%) 53.26% (0.2%) 52.97% (0.2%) 84.48% (1.0%) 87.03% (0.8%) 87.15% (1.3%)
Table 8: Comparison of different methods and coarsening levels for Flickr and PPI (transductive) datasets.
@@
2 levels 3 levels 4 levels 2 levels 3 levels 4 levels
GCN Fine grid          1 level: 86.48% (0.4%)          1 level: 74.88% (0.6%)
@@
Random and subgraphs(p=1) 86.68% (0.1%) 86.71% (0.2%) 86.77% (0.2%) 74.63% (0.3%) 75.37% (0.4%) 75.00% (0.3%)
Random and subgraphs(p=2) 85.64% (0.2%) 85.50% (0.2%) 85.84% (0.2%) 74.75% (0.2%) 74.75% (0.5%) 75.12% (0.3%)
GIN Fine grid          1 level: 84.82% (0.4%)          1 level: 73.89% (0.3%)
@@
Random and subgraphs(p=1) 84.73% (0.5%) 84.45% (0.4%) 83.86% (0.1%) 75.25% (0.5%) 75.87% (0.7%) 75.62% (1.5%)
Random and subgraphs(p=2) 84.51% (1.1%) 85.16% (0.1%) 84.48% (0.1%) 74.75% (0.7%) 74.63% (0.2%) 70.79% (19.8%)
GAT Fine grid          1 level: 85.16% (0.2%)          1 level: 72.65% (0.7%)
@@
Random and subgraphs(p=1) 85.86% (0.3%) 85.30% (0.2%) 85.47% (0.4%) 74.38% (1.3%) 72.65% (0.2%) 73.51% (0.7%)
Random and subgraphs(p=2) 85.86% (0.3%) 85.30% (0.2%) 85.47% (0.4%) 72.28% (1.6%) 72.40% (0.1%) 70.54% (1.0%)
Table 9: Comparison of different methods and coarsening levels for DBLP and Facebook datasets.
@@
2 levels 3 levels 4 levels 2 levels 3 levels 4 levels
GCN Fine grid          1 level: 75.87% (0.4%)          1 level: 78.07% (0.0%)

### v3

@@ Appendix D Cora, CiteSeer and PubMed Datasets
  Rand. & Sub.(p=1) $2.67\times 10^{2}$ 79.9% (1.4%) $8.10\times 10^{2}$ 71.2% (0.8%) $7.36\times 10^{2}$ 76.8% (0.5%)
  Rand. & Sub.(p=2) $3.46\times 10^{2}$ 80.2% (1.1%) $8.57\times 10^{2}$ 70.8% (0.4%) $9.87\times 10^{2}$ 77.5% (1.1%)
Table 9: Comparison of different coarsening methods on the Cora and CiteSeer datasets. The highest (best) performing approach in terms of test accuracy (%) is highlighted in black.
Method $\downarrow$ / Dataset $\rightarrow$ Cora CiteSeer
$L=2$ $L=3$ $L=4$ $L=2$ $L=3$ $L=4$
GCN
Fine Grid ($L=1$) 82.9% (1.1%) 69.6% (1.0%)
  Random(p=1) 81.7% (0.6%) 82.0% (0.5%) 81.7% (0.4%) 70.5% (0.3%) 70.4% (0.2%) 70.7% (0.2%)
  Random(p=2) 83.2% (0.6%) 83.0% (0.5%) 81.5% (0.3%) 66.6% (2.6%) 66.0% (0.8%) 66.2% (2.2%)
  Topk(p=1) 79.5% (0.7%) 79.9% (4.4%) 79.8% (2.2%) 68.5% (1.1%) 68.1% (0.1%) 67.4% (0.3%)
  Topk(p=2) 80.5% (2.5%) 79.6% (2.4%) 80.8% (0.8%) 66.4% (0.8%) 63.6% (2.7%) 65.8% (2.1%)
  Subgraphs(p=1) 82.7% (0.7%) 83.5% (0.4%) 81.3% (0.1%) 71.6% (0.5%) 71.2% (0.4%) 71.3% (0.4%)
  Subgraphs(p=2) 82.8% (0.4%) 83.1% (0.5%) 81.8% (0.7%) 71.5% (0.5%) 70.9% (0.2%) 70.9% (0.4%)
  Rand. & Sub.(p=1) 81.7% (0.4%) 81.5% (0.3%) 81.5% (0.3%) 71.4% (0.5%) 70.0% (0.0%) 70.7% (0.3%)
  Rand. & Sub.(p=2) 82.7% (0.3%) 82.6% (0.5%) 80.8% (0.1%) 70.8% (0.3%) 70.9% (0.2%) 70.7% (0.1%)
GIN
Fine Grid ($L=1$) 76.4% (1.6%) 63.8% (2.6%)
  Random(p=1) 76.6% (1.5%) 76.2% (2.7%) 77.6% (2.2%) 66.3% (7.0%) 66.5% (2.2%) 66.4% (1.1%)
  Random(p=2) 73.1% (0.3%) 77.2% (0.6%) 71.6% (1.6%) 60.6% (2.6%) 60.5% (0.8%) 62.8% (1.3%)
  Topk(p=1) 77.5% (1.3%) 77.6% (5.6%) 79.4% (1.9%) 67.3% (1.3%) 64.8% (3.9%) 66.5% (1.5%)
  Topk(p=2) 76.2% (1.6%) 74.6% (3.0%) 75.6% (2.3%) 61.1% (1.1%) 60.5% (2.0%) 60.0% (2.9%)
  Subgraphs(p=1) 77.5% (0.9%) 78.7% (1.1%) 78.2% (1.4%) 68.2% (0.5%) 68.8% (1.6%) 68.7% (1.4%)
  Subgraphs(p=2) 76.2% (8.57%) 76.9% (1.72%) 77.3% (8.55%) 60.1% (0.29%) 65.0% (1.8%) 66.6% (1.0%)
  Rand. & Sub.(p=1) 80.2% (2.5%) 76.9% (1.0%) 76.8% (2.1%) 66.0% (0.9%) 65.3% (1.2%) 65.5% (1.1%)
  Rand. & Sub.(p=2) 75.5% (0.7%) 75.9% (0.8%) 74.6% (22.5%) 64.1% (2.2%) 66.0% (0.2%) 66.7% (1.0%)
GAT
Fine Grid ($L=1$) 79.6% (2.05%) 67.2% (6.5%)
  Random(p=1) 80.1% (0.8%) 80.3% (0.2%) 79.8% (0.8%) 71.1% (0.4%) 72.9% (1.5%) 71.4% (0.3%)
  Random(p=2) 82.0% (0.5%) 79.5% (0.6%) 79.5% (1.9%) 71.6% (1.1%) 70.7% (1.0%) 70.6% (1.5%)
  Topk(p=1) 77.3% (1.8%) 79.8% (0.6%) 77.5% (2.9%) 67.2% (1.5%) 66.8% (0.6%) 70.8% (1.3%)
  Topk(p=2) 79.6% (2.3%) 76.8% (6.7%) 79.0% (1.6%) 67.4% (2.2%) 65.2% (0.7%) 67.0% (0.2%)
  Subgraphs(p=1) 80.2% (2.6%) 78.5% (0.8%) 79.5% (0.8%) 69.1% (0.6%) 70.2% (0.6%) 68.4% (2.7%)
  Subgraphs(p=2) 80.5% (1.5%) 80.7% (1.2%) 79.9% (0.9%) 70.1% (0.5%) 68.6% (0.3%) 66.0% (0.2%)
  Rand. & Sub.(p=1) 80.6% (0.2%) 79.9% (1.4%) 80.3% (0.9%) 70.8% (0.1%) 71.2% (0.8%) 71.2% (0.3%)
  Rand. & Sub.(p=2) 81.1% (1.1%) 80.2% (1.1%) 78.7% (0.4%) 70.3% (0.4%) 70.8% (0.4%) 71.1% (0.3%)
Table 10: Comparison of different coarsening methods on the PubMed dataset. The highest (best) performing approach in terms of test accuracy (%) is highlighted in black.
Method $\downarrow$ / Levels $\rightarrow$ $L=2$ $L=3$ $L=4$
GCN
Fine Grid ($L=1$) 77.5% (0.4%)
  Random(p=1) 77.6% (0.8%) 77.5% (0.6%) 77.5% (0.4%)
  Random(p=2) 74.2% (0.4%) 73.8% (0.4%) 74.2% (0.7%)
  Topk(p=1) 77.8% (0.9%) 78.2% (0.3%) 78.2% (0.6%)
  Topk(p=2) 75.4% (0.4%) 74.5% (0.4%) 74.2% (0.6%)
  Subgraphs(p=1) 77.7% (0.2%) 78.0% (0.3%) 78.7% (1.3%)
  Subgraphs(p=2) 74.8% (1.2%) 76.1% (0.3%) 76.5% (1.1%)
  Rand. & Sub.(p=1) 77.4% (0.5%) 77.8% (0.8%) 76.7% (0.3%)
  Rand. & Sub.(p=2) 74.1% (0.7%) 73.9% (0.4%) 75.6% (0.9%)
GIN
Fine Grid ($L=1$) 75.4% (0.4%)
  Random(p=1) 75.6% (1.7%) 75.7% (2.0%) 76.2% (1.3%)
  Random(p=2) 75.6% (2.9%) 75.1% (0.9%) 74.3% (1.2%)
  Topk(p=1) 76.6% (0.8%) 75.6% (1.6%) 75.5% (1.6%)
  Topk(p=2) 76.0% (1.8%) 76.5% (1.5%) 66.1% (6.4%)
  Subgraphs(p=1) 75.8% (1.2%) 77.2% (1.2%) 77.2% (0.7%)
  Subgraphs(p=2) 75.6% (2.8%) 75.2% (1.5%) 55.4% (4.7%)
  Rand. & Sub.(p=1) 75.2% (0.8%) 76.1% (2.8%) 75.0% (0.7%)
  Rand. & Sub.(p=2) 77.8% (5.6%) 75.9% (0.5%) 76.4% (2.4%)
GAT
Fine Grid ($L=1$) 75.5% (0.5%)
  Random(p=1) 76.5% (0.3%) 77.5% (1.5%) 76.4% (0.8%)
  Random(p=2) 75.2% (1.0%) 77.4% (1.2%) 78.6% (2.5%)
  Topk(p=1) 76.1% (0.2%) 76.1% (0.9%) 75.6% (2.5%)
  Topk(p=2) 74.2% (1.0%) 74.9% (0.7%) 73.3% (1.0%)
  Subgraphs(p=1) 77.0% (1.0%) 75.4% (4.6%) 75.8% (1.6%)
  Subgraphs(p=2) 75.1% (0.85%) 77.3% (1.1%) 78.2% (3.0%)
  Rand. & Sub.(p=1) 77.3% (0.8%) 76.8% (0.5%) 76.0% (1.9%)
  Rand. & Sub.(p=2) 75.7% (1.8%) 77.5% (1.1%) 76.4% (1.9%)
Appendix E Additional Transductive Learning datasets
We further test our methods using additional datasets. We provide results for Flickr Zeng et al. (2019) and PPI (transductive) Yang et al. (2020), showing both timing analysis, in Table 11, and full multiscale training results in Table 13. Our results show that while achieving comparable accuracy to the original fine grid, our methods show significant timing improvement both in a theoretical manner as well as in FLOPs analysis, which we calculate as explained in Section 5.
@@ Appendix E Additional Transductive Learning datasets
Table 11: Comparison of different coarsening methods on the Flickr and PPI (transductive) datasets. The lowest (best) FLOPs count for each GNN appears in blue. Time represents the training epoch time in milliseconds. The lowest (best) time for a given network and level is highlighted in green. The highest (best) performing approach in terms of test accuracy (%) is highlighted in black.
@@
FLOPs(M) Time(ms) Test Acc(%) FLOPs(M) Time(ms) Test Acc(%)
GCN
Fine grid $1.72\times 10^{4}$ 68.90 54.77% (0.2%) $8.09\times 10^{3}$ 64.66 70.74% (0.2%)
@@
Rand. & Sub. $4.44\times 10^{3}$ 49.74 55.85% (1.6%) $1.63\times 10^{3}$ 68.06 71.50% (0.0%)
GIN
Fine grid $3.58\times 10^{4}$ 76.29 52.05% (1.0%) $1.34\times 10^{4}$ 48.76 85.67% (7.2%)
@@
Rand. & Sub. $9.01\times 10^{3}$ 53.51 52.00% (1.8%) $3.16\times 10^{3}$ 68.52 90.83% (12.3%)
GAT
Fine grid $1.53\times 10^{4}$ 103.54 52.71% (0.9%) $4.57\times 10^{3}$ 192.20 82.37% (3.1%)
@@
Rand. & Sub. $3.99\times 10^{3}$ 62.38 53.26% (0.2%) $9.21\times 10^{2}$ 79.21 87.03% (0.8%)
Table 12: Datasets statistics.
Dataset Nodes Edges Features Classes
Cora Sen et al. (2008) 2,708 10,556 1,433 7
Citeseer Sen et al. (2008) 3,327 9,104 3,703 6
Pubmed Sen et al. (2008) 19,717 88,648 500 3
Flickr Zeng et al. (2019) 89,250 899,756 500 7
WikiCS Mernyei & Cangea (2020) 11,701 216,123 300 10
DBLP Bojchevski & Günnemanng (2017) 17,716 105,734 1,639 4
Facebook Yang et al. (2020) 4,039 88,234 1,283 193
BlogCatalog Yang et al. (2020) 5,196 343,486 8,189 6
PPI (transductive) Yang et al. (2020) 56,944 1,612,348 50 121
Table 13: Comparison of different coarsening methods on the Flickr and PPI (transductive) datasets. The highest (best) performing approach in terms of test accuracy (%) is highlighted in black.
@@
$L=2$ $L=3$ $L=4$ $L=2$ $L=3$ $L=4$
GCN
Fine Grid ($L=1$) 54.77% (0.2%) 70.74% (0.2%)
@@
  Rand. & Sub.s 55.65% (0.1%) 55.85% (1.6%) 55.08% (0.2%) 70.88% (0.1%) 71.50% (0.0%) 72.89% (0.1%)
GIN
Fine Grid ($L=1$) 52.05% (1.0%) 85.67% (7.2%)
@@
  Rand. & Sub. 52.74% (1.0%) 52.00% (1.8%) 52.50% (2.0%) 84.25% (4.1%) 90.83% (12.3%) 93.31% (12.1%)
GAT
Fine Grid ($L=1$) 52.71% (0.9%) 82.37% (3.1%)
@@
  Rand. & Sub. 52.70% (0.0%) 53.26% (0.2%) 52.97% (0.2%) 84.48% (1.0%) 87.03% (0.8%) 87.15% (1.3%)
Table 14: Comparison of different coarsening methods on the DBLP and Facebook datasets. The highest (best) performing approach in terms of test accuracy (%) is highlighted in black.
@@
$L=2$ $L=3$ $L=4$ $L=2$ $L=3$ $L=4$
GCN
Fine Grid ($L=1$) 86.48% (0.4%) 74.88% (0.6%)
@@
  Rand. & Sub.(p=1) 86.68% (0.1%) 86.71% (0.2%) 86.77% (0.2%) 74.63% (0.3%) 75.37% (0.4%) 75.00% (0.3%)
  Rand. & Sub.(p=2) 85.64% (0.2%) 85.50% (0.2%) 85.84% (0.2%) 74.75% (0.2%) 74.75% (0.5%) 75.12% (0.3%)
GIN
Fine Grid ($L=1$) 84.82% (0.4%) 73.89% (0.3%)
@@
  Rand. & Sub.(p=1) 84.73% (0.5%) 84.45% (0.4%) 83.86% (0.1%) 75.25% (0.5%) 75.87% (0.7%) 75.62% (1.5%)
  Rand. & Sub.(p=2) 84.51% (1.1%) 85.16% (0.1%) 84.48% (0.1%) 74.75% (0.7%) 74.63% (0.2%) 70.79% (19.8%)
GAT
Fine Grid ($L=1$) 85.16% (0.2%) 72.65% (0.7%)
@@
  Rand. & Sub.(p=1) 85.86% (0.3%) 85.30% (0.2%) 85.47% (0.4%) 74.38% (1.3%) 72.65% (0.2%) 73.51% (0.7%)
  Rand. & Sub.(p=2) 85.86% (0.3%) 85.30% (0.2%) 85.47% (0.4%) 72.28% (1.6%) 72.40% (0.1%) 70.54% (1.0%)
Table 15: Comparison of different coarsening methods on the BlogCatalog and WikiCS datasets. The highest (best) performing approach in terms of test accuracy (%) is highlighted in black.
@@
$L=2$ $L=3$ $L=4$ $L=2$ $L=3$ $L=4$
GCN
Fine Grid ($L=1$) 75.87% (0.4%) 78.07% (0.0%)